Home‎ > ‎Crew 134‎ > ‎

120-CommandersReport

Commanders Report
01/20/2014
Ashley Dale

Today was a productive and busy day. We travelled both north and south of the HAB and found two ideal sites for rover field-testing. We mounted hardware onto the 270lb rover and are in the process of getting it fully-operational for field-testing. We collected soil samples for two separate studies, extreme weather-testing of extremophiles, and hydrogen/oxygen extraction from soil hardware-testing, that are being conducted now. We conducted reconnaissance of areas of interest for ESA/OU hardware-testing.
 
Just after midnight tonight, we’ll have a film crew arriving, contracted by the Science Channel, as part of Discovery. This group will record various aspects of the work we’re doing here, starting with a tele-surgery protocol development study in collaboration with four entities around the world, including a group at the Concordia base in Antarctica.
 
There is never a dull moment here!

--------------------------------------

CommanderCheckIn Report
01/20/2014
Ashley Dale
Crew Physical Health:
Michaela has had a scratchy throat since transit to MDRS, but it 
doesn't sound serious - it's not bad nor is it getting worse (HSO is 
keeping an eye on her). Some of the crew were working very late last 
night, and by consequence are quite tired, and will be working into 
the very early hours on a tele-surgery study in coordination with 4 
other groups around the world. I need to keep a close eye on 
everyone's level of exhaustion -- we don't want anyone burning out 
until the end of our expedition.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:
EVA 1 - Successful collection of geological samples for extremophile experiments to occur over the coming days. Found great location for rover field-testing 1km south of HAB, and 500m NE of HAB. Successful recon and pictures taken of areas of interest for hardware testing ESA/OU soil sample collection/containment.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:
EVA 2 - ESA/OU hardware successfully mounted to rover. Technical issues with rover are in the process of being resolved, before fully operational for field-testing - expected EVA time required: 10hrs. Successfully completed soil sample collection for analysis of water content.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:
EVA 3 - Medical EVA postponed to tomorrow. Film crew from Science Channel are a day late, so we are saving EVA for them. Did a quick dry-run just at the beginning of EVA 1.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
   Points of interest within engineering report, but nothing out of the ordinary.
Rep Sched:
   n/a
Tomorrow Plan:
   Science Channel arriving at 01h00. Tele-surgery study to be conducted with 4 other channels 02h00-05h00. Postponed medical EVA to be conducted in afternoon. Vibha and Michaela to conduct experiments on samples collected from EVAs today.
Inventory:
We need freeze dried tomatoes/carrots/other assorted vegetables, black pepper, milk, toilet paper, pasta, green beans, dish soap, hand sanitizer, disinfectant spray, bandages, baking powder
Support Req:
   n/a
Miscelleneous:
n/a